OpenStudy (anonymous):

The circumference of a circle and the perimeter of a rectangle are equal. The radius of the circle is equal to the width of the rectangle. If the radius of the circle is 10.5 ft, what is the length of the rectangle? Use 3.14 for .

OpenStudy (anonymous):

\[2 \pi r = 2(L+r)\]

OpenStudy (anonymous):

22.47 ft is the length\[2L=2 \pi r -2r\]\[L=(\pi -1)r\]